OpenStudy (anonymous):

How do I solve this type of problem? The square root of 56 over the square root of 50?

OpenStudy (anonymous):

Break down 56 and 50 to prime numbers.

OpenStudy (anonymous):

And remember that: \[\sqrt{x \times x} = \sqrt{x} \times \sqrt{x}\]

OpenStudy (anonymous):

ok @ankh then what?

OpenStudy (anonymous):

Then you can reduce the expression, because you will have things that will repeat in the nominator and in the denominator

OpenStudy (anonymous):

Another important rule: \[\sqrt{x} \times \sqrt{x} = x\]

OpenStudy (anonymous):

ok so now I have 28 over 25 in a square root

OpenStudy (anonymous):

A radical times a radical cancels out a radical? @ankh

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

28 and 25? Make sure you do a complete prime factorization

OpenStudy (anonymous):

I got it! I did the prime thing it's\[\frac{ 2\sqrt{7} }{ 5 }\] Thank you!!!

OpenStudy (anonymous):

:)